I need to solve this linear equation, but i just do not know how. I need the proper steps to solving this problem. Please help!!!
3x-2.5y=7.125
2.5x-3y=7.3125

Answer by tutorcecilia(2152)

3x-2.5y=7.125
3x-3x-2.5y=-3x+7.125 [solve for "y" by isolating the y-variable]
-2.5y=-3x+7.125
-2.5y/-2.5=(-3x/-2.5) + (7.125/-2.5)
y=1.2x-2.85
.
2.5x-3y=7.3125
2.5x-2.5x-3y=-2.5x+7.3125
-3y=-2.5x+7.3125
-3y/-3=(-2.5x/-3)+(7.3125/-3)
y=0.8333x-2.4375